# Energy-Efficiency and Spectral-Efficiency Trade-off in Distributed Massive-MIMO Networks

### v2

@@ IV The SCA based Energy Efficiency maximization
As argued above, the optimization problem in (5) is, generally, computationally hard. Therefore, in the following we provide an approach to find a feasible solution to this problem that guarantees a predefined system-wise sum SE. The proposed approach employs various relaxations and approximations steps, as outlined and discussed below:
  1. 1.
    As both the numerator and denominator of the objective function in (5), we first reformulate the optimization problem to reduce its solution complexity, without affecting the optimality of its solution.
  2. 2.
    The MINLP problem is converted into a non-linear programming (NLP) problem by relaxing the binary constraint into continuous constraint.
  3. 3.
    The non-linear term inside the logarithm is addressed by introducing auxiliary variables and approximating the non-convex term $\Gamma_{t}$ using the quadratic transformation[15], making it tractable.
  4. 4.
    The fractional form of the objective function is converted into a tractable form using the auxiliary variables and the quadratic transformation[15].
  5. 5.
    The coupling between $\bm{\eta^{u}}$ and D introduces non-convexity, which is further handled in the same way as in [11], to ensure the overall problem becomes convex and tractable.
  6. 6.
    Finally, an iterative approach is employed to refine the approximated terms towards the original terms, to ensure an accurate solution.
Step 1: In the objective function (5a), the numerator and denominator both contain the sum SE term, so instead of maximizing objective function (5a), we maximize new objective function $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}}$. Then the new problem becomes:
$\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}$ $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}},$ (6a)
$0\leq\eta^{u}_{t}\leq 1,\forall t\in\mathcal{T},$ (6c)
$w\sum_{t=1}^{T}\log_{2}(1+\Gamma_{t})\geq\text{SE}^{\text{QoS}},\forall t\in\mathcal{T},$ (6d)
$\sum\limits_{m=1}^{M}d_{mt}\geq 1,\forall t\in\mathcal{T},$ (6e)
As maximizing the objective function (6a) is same as maximizing the original objective function (5a), thus solving Problem (6) is equivalent to solving the original Problem (5).
Step 2: We relax the binary variable $d_{mt}$ to a continuous constraint, as complexity of solving a mixed-integer type problem is exponential. Thus, constraint (6b) can be written as:
@@
Remark : In objective function (6a), the power consumption term behaves as the penalty for the AP-UE association matrix and thus forcing the $d_{mt}$ to take values either close to one or zero. Thus, maintaining the feasibility of our original constraint (6b).
Step 3: Given that $\Gamma_{t}$ is a non-linear term inside the logarithmic function, we introduce auxiliary variable $\Gamma^{*}_{t}$ as the lower bound on $\Gamma_{t}$. This substitution removes the non-linearity from the logarithmic term by replacing $\Gamma_{t}$ with $\Gamma^{*}_{t}$ in both the objective function (6a) and constraint (6d). Consequently, a new constraint is introduced in Problem (6) to ensure the feasibility of a new variable $\Gamma^{*}_{t}$. The additional constraint is given as:
@@
Still, the new constraint (8) is non-convex to fractional non-convex term, $\Gamma_{t}$. Thus, the lower bound of the right hand side of constraint (8) is introduced, by using the quadratic transformation[15], and is given by:
$\Gamma_{t}\geq 2z_{t}\sqrt{\eta^{u}_{t}p_{u}\left|\sum\limits_{m\in\mathcal{Z}_{t}}d_{mt}\gamma_{mt}+A\sum\limits_{m\in\mathcal{Q}_{t}}d_{mt}\gamma_{mt}\right|^{2}}-z_{t}^{2}I_{t},$ (9)
where the optimal value of $z_{t}$ is given by:
$z^{*}_{t}=\frac{\sqrt{\eta^{u}_{t}p_{u}\left|\sum\limits_{m\in\mathcal{Z}_{t}}d_{mt}\gamma_{mt}+A\sum\limits_{m\in\mathcal{Q}_{t}}d_{mt}\gamma_{mt}\right|^{2}}}{I_{t}}.$ (10)
Thus, constraint (8) can be written as, :
$\Gamma^{*}_{t}\leq 2z_{t}\sqrt{\eta^{u}_{t}p_{u}\left|\sum\limits_{m\in\mathcal{Z}_{t}}d_{mt}\gamma_{mt}+A\sum\limits_{m\in\mathcal{Q}_{t}}d_{mt}\gamma_{mt}\right|^{2}}-z^{2}_{t}I_{t}.$ (11)
The feasibility of constraint (11) lies in the fact that constraint (11) is the lower bound of constraint (8), as can be seen from inequality (9). Therefore, Problem (6) after incorporating the new auxiliary variable $\Gamma^{*}_{t}$ and constraint (11) becomes:
$\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}$ $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma^{*}_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}},$ (12a)
subject to : $w\sum_{t=1}^{T}\log_{2}(1+\Gamma^{*}_{t})\geq\text{SE}^{\text{QoS}},\forall t\in\mathcal{T},$ (12b)
@@
In order to solve Problem (12), constraint (11) attempts to achieve equality. The value of $z_{t}$ in (11) needs to be updated after each iteration using equation (10). After substitution of $z_{t}$ value in (11), constraint (11) becomes equivalent to constraint (8). As the iterations proceeds in our proposed iterative solution, $\Gamma^{*}_{t}$ become equal to $\Gamma_{t}$. This implies that solving Problem (12) is equivalent to solving Problem (6).
@@
$\underset{\bm{\eta^{u}},\textbf{D},u,v}{\max~{}}$ $\ \ \frac{u}{v}$ (13a)
subject to : $w\sum_{t=1}^{T}\log_{2}(1+\Gamma^{*}_{t})\geq\text{SE}^{\text{QoS}},\forall t\in\mathcal{T},$ (13b)
$wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma^{*}_{t})\geq{u},$ (13c)
@@
Solving Problem (13) is equivalent to solving (12) as constraints (13c) and (13d) must attain equality as iterations proceed. Still, we need to handle the fractional objective function (13a). The objective function (12a), which is a fraction of linear variables can also be approximated using the quadratic transformation[15] as:
@@
where the optimal value of $b$ is given by:
$b^{*}=\frac{\sqrt{wB\sum_{t=1}^{T}\text{SE}^{u}_{t}}}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}}.$ (15)
Since, inequality in (14) provides a lower bound and does not impact the feasibility as maximizing $2b\sqrt{u}-b^{2}v$ also maximizes $\frac{u}{v}$, thus the new formulation can be written as:
@@
where $\textbf{z}=\{z_{t}\}_{t\in\mathcal{T}}$.
The value of $b$ in objective function (16a) is updated using equation (15) after each iteration. As the iterations proceeds, objective function (16a) tries to achieve the values same as of objective function (13a). This indicate that solving Problem (16) is equivalent to solving Problem objective function (13).
Step 5: After fixing the parameters $b$ and z, the only non-convex term in the above formulation (16) is constraint (11) due to the coupling of $\bm{\eta^{u}}$ and D. It should be noted that, the right side of constraint (11) is bi-conacve in $\bm{\eta^{u}}$ and D. Thus, Problem (16) can be solved by an alternate optimization technique. First, we fix D, $b$ and z, then solve the following optimization problem:
@@
After solving Problem (17), we calculate the $b$ and z. Then, we fix the $\bm{\eta^{u}}$, $b$ and z, then solve the following optimization problem:
@@
Step 6: After solving Problem (18), we again calculate $b$ and z and then solve Problem (17). The iterative steps in brief are illustrated in the Algorithm 1.
@@
Algorithm 1 Proposed Algorithm
@@ V Numerical Simulations
We conduct numerical simulations over a geographic area of $1\times 1$ square kilometer with APs and UEs distributed randomly. To emulate an infinitely large network, we employ a wrap-around topology as described in [8]. The large-scale fading coefficients follow a three-slope path loss model, with shadow fading modeled by an 8 dB standard deviation. In our simulations, we fix $T=30$, $M=40$, $A=8$, $L_{c}=200$, and $L_{p}=5$. Unless explicitly mentioned otherwise, other parameters remain consistent with those in [2]. We average the numerical results across $100$ simulation runs.
Figure 2: The EE versus the sum SE threshold versus the maximum power level.
@@
Fig. 2 represents the 3-D plot between the EE, the sum SE threshold and the maximum uplink power ($p_{n}p_{u}$) that UEs can attain. In this plot, the sum SE threshold varies from $75$ to $85$ Bits/s/Hz, and the maximum uplink UE power is varied from $0.1$ to $1$ W. For the maximum power level of 0.1 W, EE decreases from $2.16\times 10^{-3}$ to $2.64\times 10^{-4}$ Gbits/Joule as the $\text{SE}^{QoS}$ increases from $75$ to $85$ Bits/s/Hz. A similar trend is observed at a power level of 0.2 W, the EE reduces from $2.5\times 10^{-3}$ to $3.42\times 10^{-4}$ Gbits/Joule for the same increase in $\text{SE}^{QoS}$. This decreasing pattern persists across higher power levels from 0.4 to 1 W, consistently showing a reduction in EE as $\text{SE}^{QoS}$ increases. This decline is attributed to the necessity of either allocating more APs to UEs, increasing the power allocated to UEs, or both, to enhance the SE. Consequently, this increases the overall power consumption and diminishes the network’s EE. We note an increase in the EE with rising maximum uplink power up to a certain point for the fixed sum SE threshold. Notably, the EE begins to saturate beyond a power level of 0.4 W for lower $\text{SE}^{QoS}$. The optimization algorithm, thereby, increases power consumption only when it contributes to the enhanced EE. For, the higher $\text{SE}^{QoS}$, however, the EE continues to increase with the rise in maximum power level, indicating that higher power levels contribute more significantly to achieving higher SE.
Figure 3: The EE versus the sum SE threshold versus number of APs for $p_{n}p_{u}=0.1$.
@@
Fig. 3 represents the 3-D plot between the EE, sum SE threshold and number of APs ($M$). For $M=40$, the EE decreases from $3.75\times 10^{-3}$ to $5.49\times 10^{-5}$ Gbits/Joule as the $\text{SE}^{QoS}$ increases from $70$ to $100$ Bits/s/Hz. A similar trend is observed at $M=80$, the EE declines from $4.53\times 10^{-3}$ to $2.8\times 10^{-3}$ Gbits/Joule for the same increase in $\text{SE}^{QoS}$. This decreasing pattern persists as the number of APs increases to 80 from 40, consistently showing a reduction in the EE as $\text{SE}^{QoS}$ rises. This consistent decrease across AP configurations indicates that higher spectral efficiency necessitate more power per unit of data transmitted, thereby reducing the EE. For the fixed $\text{SE}^{QoS}$, we observe an increase in the EE with the increase in the number of APs. This indicates that the rate of increase in the SE is more than the rate of increase in circuit power due to the increase in the number of APs. However, this increase is modest at lower $\text{SE}^{QoS}$ values such as $70$ and $75$ Bits/s/Hz. For the lower values of $M$, the optimization algorithm is easily able to achieve these lower $\text{SE}^{QoS}$ values without allocating more APs to UEs, thus not increasing the power consumption. For higher values of $M$, due to increase in the circuit power, the total power consumption increases. However, with more APs, the likelihood of having APs closer to UES increases, resulting in stronger channel condition. This proximity enhances both the SE and the EE by offseeting the impact of increases circuit power. For the higher $\text{SE}^{QoS}$ values, when the number of APs are low, in order to maintain the high sum SE, the optimization algorithm allocate more APs to UEs, thus significantly reducing the EE due to increase in power consumption. However, when the number of APs is larger for these higher $\text{SE}^{QoS}$ values, the optimizing algorithm allocates fewer APs per UEs, hence reducing the power consumption and thus enhancing the EE.
Figure 4: The EE versus the sum SE threshold versus number of UEs for $p_{n}p_{u}=0.1$.
@@
Fig. 4 represents the 3-D plot between the EE, the sum SE threshold and number of UEs ($T$). For a fixed $T$, an increase in the $\text{SE}^{QoS}$ leads to a decline in EE. At lower $\text{SE}^{QoS}$ values, the EE initially increases with $T$ but eventually decreases when $T$ becomes very high. This decline is attributed to the increased circuit power consumption required to accommodate more users at lower $\text{SE}^{QoS}$ values. Conversely, at higher $\text{SE}^{QoS}$ values, the EE consistently improves as a larger number of UEs facilitates achieving these higher $\text{SE}^{QoS}$ values.
@@ VI Conclusion and Future Work
This work has systematically explored the complex relationships among energy efficiency, spectral efficiency, and resource allocation parameters such as the number of APs and uplink power levels in distributed massive-MIMO (D-mMIMO) systems. A consistent trade-off between EE and SE is evident across different system configurations and parameters. As the SE demands increase, the EE generally decreases, indicating that higher data transmission rates necessitate greater power expenditure. This work also indicates the role of power allocation and the AP-UE association to achieve desired SE, while maintaining the satisfactory EE. The insights from this study underscore the necessity for sophisticated resource allocation strategies that prioritize both EE and SE, which is especially crucial in the era of 5G and 5G Advanced technologies. These technologies, generally, enhance SE but increase power consumption as well, making the study of this trade-off essential to maintain the environmental and the economic sustainability.
A compelling future direction would involve drawing detailed comparisons of the EE-SE trade-offs across different network architectures, including small-cell networks, traditional mMIMO and the D-mMIMO. Additionally, the influence of modulation and coding schemes, which is not considered in this study, presents an interesting direction for further exploration.

### v3

@@ IV The SCA based Energy Efficiency maximization
We provide an approach to find a feasible solution to Problem (1). The proposed approach employs various relaxations and approximations steps, as outlined and discussed below:
Step 1: In the objective function (1a), the numerator and denominator both contain the sum SE term. So, we reformulate the optimization problem to reduce its solution complexity by maximizing a new objective function $\frac{B\,\text{SE}^{u}|_{\Gamma_{t}}}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}}$. It should be noted that maximizing the new objective function is same as maximizing the original objective function in (1a), thus preserving the optimality of the solution.
Step 2: We relax the binary variable $d_{mt}$ to a continuous one to convert the given MINLP problem to a non-linear programming (NLP) problem, to write constraint (1b) as:
@@
Remark : In the new objective function, the power consumption term behaves as the penalty for the AP-UE association matrix, forcing the $d_{mt}$ to take values either close to one or zero. Thus, maintaining the feasibility of our original constraint (1b).
Step 3: Given that $\Gamma_{t}$ is non-linear, we introduce an auxiliary variable $\Gamma^{*}_{t}$ as the lower bound on $\Gamma_{t}$. This substitution removes the non-linearity from the logarithmic term by replacing $\Gamma_{t}$ with $\Gamma^{*}_{t}$ in both the new objective function and constraint (1d). Consequently, a new constraint is introduced in the reformulated problem to ensure the feasibility of the new variable $\Gamma^{*}_{t}$. The additional constraint is given as:
@@
However, this constraint is non-convex due to the fractional non-convex term, $\Gamma_{t}$. Thus, to make it tractable, we use the quadratic transformation[15], and replace constraint (3) by:
$\Gamma^{*}_{t}\leq 2z_{t}\sqrt{\text{DS}_{t}}-z^{2}_{t}I_{t}.$ (4)
The feasibility of constraint (4) lies in the fact that $\Gamma_{t}\geq 2z_{t}\sqrt{\text{DS}_{t}}-z_{t}^{2}I_{t}$. Thus, satisfying constraint (4) implies the satisfaction of constraint (3). Also, constraint (4) is marginally concave in $z_{t}$ and the optimal value of $z_{t}$ is given by $z^{*}_{t}=\sqrt{\text{DS}_{t}}/{I_{t}}$. Therefore, Problem (1) after incorporating the new auxiliary variable $\Gamma^{*}_{t}$ and constraint (4) becomes:
$\underset{\bm{\eta^{u}},\textbf{D},\textbf{z},\bm{\Gamma^{*}}}{\max~{}}$ $\frac{B\,\text{SE}^{u}|_{\Gamma^{*}_{t}}}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}},$ (5a)
subject to: $\text{SE}^{u}|_{\Gamma^{*}_{t}}\geq\text{SE}^{\text{QoS}},\forall t\in\mathcal{T},$ (5b)
@@
where $\textbf{z}=\{z_{t}\}_{t\in\mathcal{T}}$ and $\bm{\Gamma^{*}}=\{\Gamma^{*}_{t}\}_{t\in\mathcal{T}}$. In order to solve Problem (5), constraint (4) attempts to achieve equality. The value of $z_{t}$ in (4) needs to be updated after each iteration by its optimal value, $z^{*}_{t}$. After substitution of $z_{t}$ value in (4), constraint (4) becomes equivalent to constraint (3). As the iterations proceeds in our proposed iterative solution, $\Gamma^{*}_{t}$ become equal to $\Gamma_{t}$. This implies that solving Problem (5) is equivalent to solving Problem (1).
@@
$\underset{\bm{\eta^{u}},\textbf{D},\textbf{z},\bm{\Gamma^{*}},u,v}{\max~{}}$ $\ \ \frac{u}{v},$ (6a)
subject to: $\text{SE}^{u}|_{\Gamma^{*}_{t}}\geq\text{SE}^{\text{QoS}},\forall t\in\mathcal{T},$ (6b)
$B\,\text{SE}^{u}|_{\Gamma^{*}_{t}}\geq{u},$ (6c)
@@
Solving Problem (6) is equivalent to solving (5) as constraints (6c) and (6d) must attain equality as iterations proceed. Still, we need to handle the fractional objective function (6a). The objective function (5a), which is a ratio of linear terms can be approximated using the quadratic transformation[15] as:
@@
Since, $2b\sqrt{u}-b^{2}v$ is the lower bound to $\frac{u}{v}$, and thus maximizing $2b\sqrt{u}-b^{2}v$ also maximizes $\frac{u}{v}$. Therefore this transformation does not impact the feasibility of objective function (6a). Also, $2b\sqrt{u}-b^{2}v$ is marginally concave in $b$, and thus the optimal value of $b$ is is given by:
$b^{*}=\frac{\sqrt{B\,\text{SE}^{u}|_{\Gamma^{*}_{t}}}}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}}.$ (8)
The new formulation can be rewritten as:
@@
The value of $b$ in objective function (9a) is updated using equation (8) after each iteration. As the iterations proceed, the objective function (9a) tries to achieve the same values as the objective function (6a). This establishes that solving Problem (9) is equivalent to solving Problem (6).
Step 5: After fixing the parameters $b$ and z, the only non-convex term in the above formulation (9) is constraint (4) due to the coupling of $\bm{\eta^{u}}$ and D, which is further handled in the same way as in [11], to ensure the overall problem becomes convex and tractable. It should be noted that, the right side of constraint (4) is bi-concave in $\bm{\eta^{u}}$ and D. Thus, Problem (9) can be solved by an alternate optimization technique. First, we fix D, $b$ and z, and solve:
@@
After solving Problem (10), we calculate the $b$ and z. Then, we fix the $\bm{\eta^{u}}$, $b$ and z, then solve:
@@
Step 6: After solving Problem (11), we again calculate $b$ and z and then solve Problem (10). The exact iterative steps are captured Algorithm 1.
@@
Algorithm 1 The Proposed Algorithm
@@ V Numerical Simulations
We conduct numerical simulations over a geographic area of $1\times 1$ square kilometer with APs and UEs distributed randomly. To emulate an infinitely large network, we employ a wrap-around topology [8]. The LSFCs follow a three-slope path loss model, with shadow fading modeled by an 8 dB standard deviation. In simulations, we fix $T=30$, $M=40$, $A=8$, $L_{c}=200$, and $L_{p}=5$. Unless mentioned otherwise, other parameters remain consistent with those in [2]. We average the numerical results across $100$ simulation runs.
Figure 2: The EE versus the sum SE threshold versus the maximum power.
@@
Fig. 2 represents the 3-D plot among the EE, the sum SE threshold, and the maximum uplink power ($p_{n}p_{u}$) that UEs can attain. In this plot, the sum SE threshold varies from $75$ to $85$ Bits/s/Hz, and the maximum uplink UE power is varied from $0.1$ to $1$ W. For the maximum power level of 0.1 W, EE decreases from $2.16\times 10^{-3}$ to $2.64\times 10^{-4}$ Gbits/Joule as the $\text{SE}^{QoS}$ increases from $75$ to $85$ Bits/s/Hz. A similar trend is observed at the power level of 0.2 W, the EE reduces from $2.5\times 10^{-3}$ to $3.42\times 10^{-4}$ Gbits/Joule for the same increase in the $\text{SE}^{QoS}$. This decreasing pattern persists across higher power levels from 0.4 to 1 W, consistently showing a reduction in EE as $\text{SE}^{QoS}$ increases. This decline is attributed to the necessity of either allocating more APs to UEs, increasing the power allocated to UEs, or both, to enhance the SE. Consequently, this increases the overall power consumption and diminishes the network’s EE. We note an increase in the EE with rising maximum uplink power up to a certain point for the fixed sum SE threshold. Notably, the EE begins to saturate beyond the power level of 0.4 W for lower $\text{SE}^{QoS}$. The optimization algorithm, thereby, increases power consumption only when it contributes to the enhanced EE. For, the higher $\text{SE}^{QoS}$, however, the EE continues to increase with the rise in maximum power level, indicating that higher power levels contribute more significantly to achieving higher SE.
Figure 3: The EE versus the sum SE threshold versus the number of APs for $p_{n}p_{u}=0.1$.
@@
Fig. 3 represents the 3-D plot among the EE, sum SE threshold, and number of APs ($M$). For $M=40$, the EE decreases from $3.75\times 10^{-3}$ to $5.49\times 10^{-5}$ Gbits/Joule as the $\text{SE}^{QoS}$ increases from $70$ to $100$ Bits/s/Hz. A similar trend is observed at $M=80$, the EE declines from $4.53\times 10^{-3}$ to $2.8\times 10^{-3}$ Gbits/Joule for the same increase in $\text{SE}^{QoS}$. This decreasing pattern persists as the number of APs increases from 40 to 80, consistently showing a reduction in the EE as $\text{SE}^{QoS}$ rises. This consistent decrease across AP configurations indicates that higher spectral efficiency necessitates more power per unit of data transmitted, thereby reducing the EE. For the fixed $\text{SE}^{QoS}$, we observe an increase in the EE with the increase in the number of APs. This indicates that the rate of increase in the SE is more than the rate of increase in circuit power due to the increase in the number of APs. However, this increase is modest at lower $\text{SE}^{QoS}$ values such as $70$ and $75$ Bits/s/Hz. For the lower values of $M$, the optimization algorithm is easily able to achieve these lower $\text{SE}^{QoS}$ values without allocating more APs to UEs, thus not increasing the power consumption. For higher values of $M$, due to increase in the circuit power, the total power consumption increases. However, with more APs, the likelihood of having APs closer to UEs increases, resulting in stronger channel condition. This proximity enhances both the SE and the EE by offsetting the impact of increased circuit power. For the higher $\text{SE}^{QoS}$ values, when the number of APs are low, in order to maintain the high sum SE, the optimization algorithm allocates more APs to UEs, thus significantly reducing the EE due to increase in power consumption. However, when the number of APs is larger for these higher $\text{SE}^{QoS}$ values, the optimizing algorithm allocates fewer APs per UEs, hence reducing the power consumption and thus enhancing the EE.
Figure 4: The EE versus the sum SE threshold versus the number of UEs for $p_{n}p_{u}=0.1$.
@@
Fig. 4 shows the 3-D plot among the EE, the sum SE threshold, and the number of UEs ($T$). For a fixed $T$, an increase in the $\text{SE}^{QoS}$ leads to a decline in EE. At lower $\text{SE}^{QoS}$ values, the EE initially increases with $T$ but eventually decreases when $T$ becomes very high. This decline is attributed to the increased circuit power consumption required to accommodate more users at lower $\text{SE}^{QoS}$ values. Conversely, at higher $\text{SE}^{QoS}$ values, the EE consistently improves as a larger number of UEs facilitates achieving these higher $\text{SE}^{QoS}$ values.
@@ VI Conclusion and Future Work
This work has systematically explored the complex relationships among energy efficiency, spectral efficiency, and resource allocation parameters such as the number of APs and uplink power levels in distributed massive-MIMO (D-mMIMO) systems. A consistent trade-off between EE and SE is evident across different system configurations and parameters. As the SE demand increases, the EE generally decreases, indicating that higher data transmission rates necessitate greater power expenditure. This work also indicates the role of power allocation and the AP-UE association to achieve desired SE, while maintaining the satisfactory EE. The insights from this study underscore the necessity for sophisticated resource allocation strategies that prioritize both EE and SE, which is especially crucial in the era of 5G and 5G Advanced technologies. These technologies, generally, enhance SE but increase power consumption as well, making the study of this trade-off essential to maintain the environmental and the economic sustainability.
A compelling future direction involves drawing detailed comparisons of the EE-SE trade-offs across different network architectures, including small-cell networks, traditional mMIMO and the D-mMIMO. Additionally, the influence of modulation and coding schemes, which is not considered in this study, presents another direction for exploration.